OpenStudy (anonymous):

A regular octagon has side length 10.9 in. The perimeter of the octagon is 87.2 in and the area is 392.4 in2. A second octagon has side lengths equal to 5.45 in. Find the area of the second octagon. 61.60 in2 75.24 in2 37.06 in2 98.09 in2

OpenStudy (cwrw238):

the ratio of areas of the octagon's = ratio of squares of the sides using the sides of 2 octagons: ratio is 10.9 / 5.45 = 2/1 so the areas are in ratio 4:1 4/ 1 = 392.4 / a where a is area of the second octagon

OpenStudy (cwrw238):

so area = 392.4 / 4